40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 231     Entry time: Thu Jan 10 00:12:01 2008
Author: tobin 
Type: Summary 
Category: Locking 
Subject: DR 
[John, Tobin, Rana]

1. We found SUS_BS_SENSOR_UL to have a ratty signal and low DC value. Twiddling the cables at the BS satellite amplifier and vacuum feedthrough brought the signal back (to 0.667V), but it is still spiky, spiking up to a couple times per second. Rana suggested that these spikes might be scattered YAG laser light (as hypothesized in August). The spikes go away when we misalign the PRM or either ITM, and when we unlock the mode cleaner, lending credance to this theory. SUS_BS_SENSOR_UR also spikes, but much less frequently. We turned off C1:SUS-BS_ULSEN_SW2 and continued.

2. After dither alignment the oplev beams were centred and we were able to lock DRM plus either arm reliably (however locking in this state broke ./drstep_bang at the first ``Going DD''). We ran scripts/DRFPMI/bang/nospring/drdown_bang and were subsequently able to lock DRFPMI (i.e., full IFO) a couple times.

3. To do: Debug ./drstep_bang with just the DRM (no arms).
ELOG V3.1.3-